OpenStudy (anonymous):

-3/4q = 12 solve for q

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

9?

sammixboo (sammixboo):

Multiply both sides of the equation by -4/3 \(\dfrac{-4}{3}\times\dfrac{-3}{4}q = 12\times\dfrac{-4}{3}\) \(q = 12\times\dfrac{-4}{3}\)

sammixboo (sammixboo):

so what is 12*-4/3

OpenStudy (anonymous):

-16?

sammixboo (sammixboo):

Yup!

sammixboo (sammixboo):

q = -16

OpenStudy (anonymous):

ANOTHER?

sammixboo (sammixboo):

Sure!